x+3y=9
x+y=1 combine llike terms 4y=9
divide by 4 y=2.25
substitute x+2.25=1
i think that's wrong help please

Answer by scott8148(6628)

subtracting equations ___ x + 3y - x - y = 9 - 1 ___ 2y = 8

solve for y and substitute back to find x